# Dense Conjugacy Classes and Stability of Locally Finite Graphs

### v1

@@ 1 Introduction
A topological group with a dense conjugacy class has the Rokhlin property. Examples include the group of homeomorphisms of Cantor space and that of the Hilbert Cube [GW], as well as the symmetric group of the natural numbers [M]. For more examples, history, and context, see Kechris and Rosendal [KR]. A complete classification of which mapping class groups of connected orientable surfaces have a dense conjugacy class was obtained by Lanier and Vlamis [LV] and Hernández Hernández, Hrušák, Morales, Randecker, Sedano, and Valdez [HHMRSV]: the mapping class group of a connected orientable $2$-manifold has the Rokhlin property if and only if the manifold is either the $2$-sphere or a non-compact manifold whose genus is either zero or infinite and whose end space is self-similar with a unique maximal end.
@@
When $X$ is not self-similar, the situation for locally finite graphs becomes more complicated than for surfaces, and a full classification is currently out of reach. For example, one obstruction to having a dense conjugacy class in the setting of surfaces is the existence of a non-displaceable subsurface. There is no meaningful analogue of a non-displaceable subsurface in the setting of locally finite graphs. Other obstructions, such as containing a subgroup that is proper, open, and normal, do apply to this setting. We show that such subgroups exist in a broad variety of settings by constructing non-trivial group homomorphisms to discrete groups including $\mathbb{Z}$, $\text{Out}(F_{n})$, and the symmetric group on $n$ letters. The homomorphisms from $\text{Maps}(X)$ to $\mathbb{Z}$, which we call flux maps, generalize the flux maps defined by [DHK], which in turn are analogous to the flux maps of [MRlong, Lemma 6.7].
The flux maps of [DHK] count the number of loops passing from one end to another and are defined on the subgroup of $\text{Maps}(X)$ consisting of mapping classes which induce the identity on the end space of $X$. We extend the flux map construction to apply to all of $\text{Maps}(X)$ and to count ends in addition to loops. Intuitively, the ends we count must satisfy a maximality condition, which we call being a greatest common divisor, or gcd, to ensure that the flux maps will be finite-valued. See Definitions 5.17 and 5.27 for the precise definitions of gcds and flux maps respectively, and see Section 2.3 for a discussion of end types and the order on ends.
@@
    There exist two maximal ends with a gcd which is not of Cantor type.
@@
    There exists an end type $E(\mu)$ with $0<|E(\mu)|<\infty$.
@@
The methods we employ to prove that Theorem 1.2 (1) implies the conclusion apply to many subgroups of $\text{Maps}(X)$, as well; see Theorem 5.32 for the general statement. In particular, the following corollary follows immediately from Theorem 5.32.
@@
Any graph $X\vee C$ as in Theorem 1.4 is an example of a graph such that the mapping class group of the corresponding surface does not have a dense conjugacy class. This is an uncountable class of graphs, because it follows from [MS] and the constructions in Section 3.1 that such graphs are naturally indexed by countable ordinal numbers.
To facilitate our investigation of dense conjugacy classes, we develop a combinatorial notation for locally finite graphs based on their end space, called a signature. The definition is technical, so we instead refer to Figure 2 for examples and refer readers to Definition 3.1. We then develop the stronger notion of an ordered signature, which explicitly encodes the order on ends and is equivalent to the graph being stable. Stability (Definition 2.23) is a natural condition on the end space of a locally finite graph that ensures that it is “well-behaved.” First introduced by Mann and Rafi [MRlong], many papers have restricted to the class of stable locally finite graphs or surfaces, including works of Fanoni, Ghaswala, and McLeay [FGM] and works of Bar-Natan and Verberne [BV], along with many others.
@@
The locally finite graphs for which it is still unknown whether the mapping class has a dense conjugacy class or not roughly fall into three categories. The first category is comprised of graphs which are essentially too simple to have any predictive structure, such as $1\vee o(1)$ or $(\omega+1)\vee C\vee o(1)$. In these graphs, there do not exist ends (or loops) which are dominated by two maximal ends. The second category includes graphs such as $\{\omega^{n}+1\}_{n}\to(1\vee C)$. These are graphs where for any pair of maximal end types, there is an increasing sequence of end types dominated by those maximal end types. The third category are unstable graphs where the methods involving stable maximal ends do not apply. This includes the graphs constructed in [MRshort] and the graph in Example 4.22.
@@ Acknowledgments
Thank you to the author’s advisor Carolyn R. Abbott for her help and support throughout the project. Thank you to Ferrán Valdez for suggesting the problem and for his advice, and thank you to Nicholas G. Vlamis for many helpful conversations and pointing the author in the direction of flux maps. Thank you to George Domat for many helpful brainstorming sessions, Robbie Lyman for help visualizing intricate proper homotopy equivalences of graphs, and Kasra Rafi and Assaf Bar-Natan for conversations about complex end spaces. Thank you to Brian Udall for pointing out errors in initial drafts of Proposition 5.18 and for his suggestions, and thank you to Arya Vadnere for help in proving $3\Rightarrow 4$ in Proposition 4.4. The author was partially supported by NSF grants DMS-2106906 and DMS-2340341.
@@ 2.1 Locally finite graphs
We give an overview of the definitions and properties of locally finite graphs necessary for this paper. For further details see [AK-B] and [MRlong]. All graphs $G$ are assumed to be locally finite and connected, and $\mathcal{V}(G)$ denotes the set of vertices of $G$.
@@
The end space $\partial X_{g}\subseteq\partial X$ is the subspace of ends “accumulated by genus,” or such that every neighborhood in $X\cup\partial X$ of the end has positive genus. The set $\partial X_{g}$ is always closed in $\partial X$ and is the empty set precisely when $g(X)<\infty$ [AK-B, Section 2]. A locally finite graph is finite-type if $|\partial X|$ and $g(X)$ are both finite, and is infinite-type otherwise.
@@
We further endow $\text{Maps}(X)$ with the topology constructed in [AK-B], which we briefly describe. Let $K$ be a finite subgraph of $X$. Let $U_{K}$ be the set of mapping classes $[f]$ with a representative $f$ such that the following four conditions hold.
@@
The topology on $\text{Maps}(X)$ is the coarsest topology such that all subsets of the form $U_{K}$ are open and such that multiplication and inversion are continuous. The following two propositions summarize results in [AK-B] which are relevant to our discussion.
Proposition 2.6 ([AK-B]).
@@ 2.3 Order on ends
Ends in a locally finite graph can be endowed with a preorder [MRlong]. We state the definition and then give the relevant background.
@@
We next define stability, which is a natural condition that rules out many pathological locally finite graphs, defined in [MRlong, Section 4].
@@
    If $X$ is a tree with countable end space, then $X$ is stable [FGM, Remark 5.5].
@@
    Non-stable locally finite graphs are more complicated and it will be easier to give examples after Definition 3.1. Intuitively, if there is a sequence of ends $\{\nu_{n}\}_{n}$ converging to an end $\mu$, none of which are dominated by $\mu$ and which are all of different types, then $\mu$ must not be stable. For concrete examples, see Example 4.22 and the first bullet point under part three in Definition 4.23. There are other ways to violate stability as well; the graphs in [MRshort, Theorem 1.1] are not stable and contain a unique maximal end type.
@@
Mann and Rafi [MRlong, Section 4] show that, without loss of generality, one can take $n=2$.
@@ 3.1 Signatures
By construction, signatures are collections of symbols associated to particular locally finite graphs which are invariant under proper homotopy. Throughout the paper, we sometimes mention signatures and actually mean a particular graph with that signature; for example, we may refer to the graph $(\omega^{3}+1)\vee C$. Also note that a proper homotopy equivalence of graphs induces a homeomorphism of their end spaces [AK-B, Section 2]. Thus, for example, by an abuse of notation, we may refer to the end space of any graph with signature $(\omega^{3}+1)\vee C$ by $\partial((\omega^{3}+1)\vee C)$.
In the context of $(\{Y_{n}\}_{n},a)\to\mathcal{C}$, where each of the $Y_{n}$ are arbitrary locally finite trees, $\mathcal{C}$ is an infinite binary splitting tree, and $a$ is a specific bijection, a notation that is similar in flavor capturing behavior of the end space was developed by [Ket, Definition 0.4].
@@ 3.2 Building blocks
We now turn our attention to proving lemmas which will be useful throughout the paper. The following lemma is an implication of [MRlong, Lemma 4.17], and we include a proof for completeness.
@@
First assume that $V$ is contained within $U$. It follows by taking $x=y$ in [MRlong, Lemma 4.17] that all sufficiently small neighborhoods of $\nu$ are homeomorphic to $U$. In particular, if $W$ is a sufficiently small neighborhood of $\nu$, then there exists a homeomorphism $f:U\rightarrow W$ fixing $\nu$ and a homeomorphic copy $f(V)$ of $V$ in $W$. The set $f(V)$, being a subset of $W$, is also a sufficiently small neighborhood of $\nu$, and is therefore homeomorphic to $U$. Thus $U$ is homeomorphic to $V$. For the moreover statement, suppose that $\nu^{\prime}\sim\nu$. By [MRlong, Lemma 4.17], all sufficiently small neighborhoods of $\nu^{\prime}$ are homeomorphic to $U$ via a homeomorphism taking $\nu$ to $\nu^{\prime}$, which immediately gives the result. ∎
@@
We first prove the statement when $A=\{\nu\}$ is a singleton. Without loss of generality, assume that $X$ is in the standard form of [AK-B, Section 2]. Let $x_{0}$ be a vertex in $X$, and consider a non-backtracking path based at $x_{0}$ in $X$ limiting to $\nu$ with vertices $\{x_{n}\}_{n}$. Because $X$ is in standard form, $X\cong\{Y_{n}\}_{n}\rightarrow 1$, where each $Y_{n}$ is attached to the path limiting to $\nu$ via $x_{n}$. For each $n$, the set $\{\nu\}\cup(\bigcup_{i\geq n}\partial Y_{i})$ is clopen in $\partial X$. If, for contradiction, $V$ intersected infinitely many of these sets, then $\nu$ would be a limit point of $V$, contradicting that $V$ is closed. Thus take $U$ to be one of these sets for suitably large $n$.
@@
The end space $\partial X$ is compact and Hausdorff. By [AK-B, Definition 2.5], it embeds into Cantor space, and therefore has a countable basis consisting of clopen sets. If $\partial X$ is not homeomorphic to Cantor space, then by Brouwer’s characterization of Cantor space, it has an isolated point. Hence $X$ has a discrete end. ∎
@@
Using the construction in [AK-B, Definition 2.5], we fix an embedding of $\partial X$ into $\mathbb{R}$. For any clopen subset of the end space $U\subseteq\partial X$, there exist finitely many disjoint open intervals $I_{k}$ in $\mathbb{R}$ such that $\partial X\cap(\bigcup_{k}I_{k})=U$. Indeed, by Lemma 3.6 and compactness of $U$, there are finitely many open intervals in $\mathbb{R}$ covering $U$ with $\partial X\cap(\bigcup_{k}I_{k})=U$, and we may assume that they are disjoint because if two overlapped, we could combine them into one open interval. Thus for each $m$ we obtain finitely many disjoint open intervals $I_{m,1},...,I_{m,k_{m}}$ covering $U_{m}^{\prime}$ that are disjoint from $\partial X\setminus U_{m}^{\prime}$. $\{I_{m,i}\,:\,m,i\in\mathbb{Z}_{>0}$ and $1\leq i\leq k_{m}\}$ is a set of pairwise disjoint open intervals in $\mathbb{R}$, and thus we assume that they are indexed in an order-preserving way by $\mathbb{Z}$. Define $U_{m}:=I_{m}\cap\partial X$ so that $\{U_{m}\}_{m\in\mathbb{Z}}$ is a clopen decomposition of $\partial X\setminus V$.
@@ 4.1 Classification of local structures
Proposition 4.4 suggests that stability is a local version of self-similarity. Indeed, the bi-implication $1\Leftrightarrow 2$ is saying that a neighborhood of an end is stable if and only if a locally finite graph with end space equal to that neighborhood has self-similar end space. There is a construction in [MRshort] of graphs with non-stable end spaces which contain a unique maximal end type such that this end type is of size one or infinity. Thus, the requirement in statement $3$ that maximal ends are stable is crucial. The graphs in [MRshort] are not self-similar, and do not satisfy any of the conditions in Proposition 4.4.
@@ 4.2 The poset of local structures
Suppose for contradiction that the $\partial Z_{i}$ form a strictly decreasing sequence of local structures. We will inductively define a descending sequence $\{U_{i}\}_{i}$ of subsets of $\mathbb{R}$. We can homeomorphically map $\partial Z_{0}$ to some subset $U_{0}\subset\mathbb{R}$ of the real numbers by the construction in [AK-B, Definition 2.5], and moreover, we may assume that the diameter of $U_{0}$, which we refer to as $\text{diam}(U_{0})$, is less than one in $\mathbb{R}$. Now suppose we have defined $U_{0},...,U_{i-1}$. Because $\partial Z_{i}\prec\partial Z_{i-1}$, there exists $\nu_{i}\in U_{i-1}$ such that $\nu_{i}\sim\partial Z_{i}$. We can realize a homeomorphic copy of $\partial Z_{i}$ as a subset $U_{i}$ of $U_{i-1}$ which is clopen in $U_{i-1}$ and which contains $\nu_{i}$. Furthermore, by stability, we may assume that $\text{diam}(U_{i})\leq\frac{1}{i+1}$. Thus, the $U_{i}$ are descending, and they are all compact and non-empty, with diameter approaching zero. This implies that $\bigcap_{i}U_{i}$ consists of a single point $x\in\mathbb{R}$. Let $\nu$ be the end in $\partial Z_{0}$ which maps to $x$, and let $\{V_{i}\}_{i}$ be a descending sequence of clopen subsets of $\partial Z_{0}$ such that $V_{i}\cong U_{i}$ for all $i\in\mathbb{Z}_{\geq 0}$. Note that as $Z_{0}$ is stable, the end $\nu$ must be stable too.

### v2

@@ 1 Introduction
A topological group with a dense conjugacy class has the Rokhlin property. Examples include the group of homeomorphisms of Cantor space and that of the Hilbert Cube [GW01], as well as the symmetric group of the natural numbers [MAC86]. For more examples, history, and context, see Kechris and Rosendal [KR07]. A complete classification of which mapping class groups of connected orientable surfaces have a dense conjugacy class was obtained by Lanier and Vlamis [LV22] and Hernández Hernández, Hrušák, Morales, Randecker, Sedano, and Valdez [HHM+22]: the mapping class group of a connected orientable $2$-manifold has the Rokhlin property if and only if the manifold is either the $2$-sphere or a non-compact manifold whose genus is either zero or infinite and whose end space is self-similar with a unique maximal end.
@@
When $X$ is not self-similar, the situation for locally finite graphs becomes more complicated than for surfaces, and a full classification is currently out of reach. For example, one obstruction to having a dense conjugacy class in the setting of surfaces is the existence of a non-displaceable subsurface. There is no meaningful analogue of a non-displaceable subsurface in the setting of locally finite graphs. Other obstructions, such as containing a subgroup that is proper, open, and normal, do apply to this setting. We show that such subgroups exist in a broad variety of settings by constructing non-trivial group homomorphisms to discrete groups including $\mathbb{Z}$, $\text{Out}(F_{n})$, and the symmetric group on $n$ letters. The homomorphisms from $\text{Maps}(X)$ to $\mathbb{Z}$, which we call flux maps, generalize the flux maps defined by [DHK23], which in turn are analogous to the flux maps of [MR23, Lemma 6.7].
The flux maps of [DHK23] count the number of loops passing from one end to another and are defined on the subgroup of $\text{Maps}(X)$ consisting of mapping classes which induce the identity on the end space of $X$. We extend the flux map construction to apply to all of $\text{Maps}(X)$ and to count ends in addition to loops. Intuitively, the ends we count must satisfy a maximality condition, which we call being a greatest common divisor, or gcd, to ensure that the flux maps will be finite-valued. See Definitions 5.17 and 5.27 for the precise definitions of gcds and flux maps respectively, and see Section 2.3 for a discussion of end types and the order on ends.
@@
    All maximal ends are stable and there exist two maximal ends with a gcd which is not of Cantor type.
@@
    There exists an end type $E(\mu)$ with $1<|E(\mu)|<\infty$.
@@
The methods we employ to prove that Theorem 1.2 (1) implies the conclusion generalize to a much less restrictive class of locally finite graphs $X$, as well as many subgroups of $\text{Maps}(X)$; see Theorem 5.34 for the general statement. In particular, the following corollary follows immediately from Theorem 5.34.
@@
Any graph $X\vee C$ as in Theorem 1.4 is an example of a graph such that the mapping class group of the corresponding surface does not have a dense conjugacy class. This is an uncountable class of graphs, because it follows from [MS20] and the constructions in Section 3.1 that such graphs are naturally indexed by countable ordinal numbers.
To facilitate our investigation of dense conjugacy classes, we develop a combinatorial notation for locally finite graphs based on their end space, called a signature. The definition is technical, so we instead refer to Figure 2 for examples and refer readers to Definition 3.1. We then develop the stronger notion of an ordered signature, which explicitly encodes the order on ends and is equivalent to the graph being stable. Stability (Definition 2.23) is a natural condition on the end space of a locally finite graph that ensures that it is “well-behaved.” First introduced by Mann and Rafi [MR23], many papers have restricted to the class of stable locally finite graphs or surfaces, including works of Fanoni, Ghaswala, and McLeay [FGM21] and works of Bar-Natan and Verberne [BV23], along with many others.
@@
The locally finite graphs for which it is still unknown whether the mapping class has a dense conjugacy class or not roughly fall into three categories. The first category is comprised of graphs which are essentially too simple to have any predictive structure, such as $1\vee o(1)$ or $(\omega+1)\vee C\vee o(1)$. In these graphs, there do not exist ends (or loops) which are dominated by two maximal ends. The second category includes graphs such as $\{\omega^{n}+1\}_{n}\to(1\vee C)$. These are graphs where for any pair of maximal end types, there is an increasing sequence of end types dominated by those maximal end types. The third category are unstable graphs where the methods involving stable maximal ends do not apply. This includes the graphs constructed in [MR24] and the graph in Example 4.22.
@@ Acknowledgments
Thank you to the author’s advisor Carolyn R. Abbott for her help and support throughout the project. Thank you to Ferrán Valdez for suggesting the problem and for his advice, and thank you to Nicholas G. Vlamis for many helpful conversations and pointing the author in the direction of flux maps. Thank you to George Domat for many helpful brainstorming sessions, Robbie Lyman for help visualizing intricate proper homotopy equivalences of graphs, and Kasra Rafi and Assaf Bar-Natan for conversations about complex end spaces. Thank you to Brian Udall for pointing out multiple errors in initial drafts of Proposition 5.18 and for his suggestions, and thank you to Arya Vadnere for help in proving $3\Rightarrow 4$ in Proposition 4.4. The author was partially supported by NSF grants DMS-2106906 and DMS-2340341.
@@ 2.1 Locally finite graphs
We give an overview of the definitions and properties of locally finite graphs necessary for this paper. For further details see [AB02] and [MR23]. All graphs $G$ are assumed to be locally finite and connected, and $\mathcal{V}(G)$ denotes the set of vertices of $G$.
@@
The end space $\partial X_{g}\subseteq\partial X$ is the subspace of ends “accumulated by genus,” or such that every neighborhood in $X\cup\partial X$ of the end has positive genus. The set $\partial X_{g}$ is always closed in $\partial X$ and is the empty set precisely when $g(X)<\infty$ [AB02, Section 2]. A locally finite graph is finite-type if $|\partial X|$ and $g(X)$ are both finite, and is infinite-type otherwise.
@@
We further endow $\text{Maps}(X)$ with the topology constructed in [AB02], which we briefly describe. Let $K$ be a finite subgraph of $X$. Let $U_{K}$ be the set of mapping classes $[f]$ with a representative $f$ such that the following four conditions hold.
@@
The topology on $\text{Maps}(X)$ is the coarsest topology such that all subsets of the form $U_{K}$ are open and such that multiplication and inversion are continuous. The following two propositions summarize results in [AB02] which are relevant to our discussion.
Proposition 2.6 ([AB02]).
@@ 2.3 Order on ends
Ends in a locally finite graph can be endowed with a preorder [MR23]. We state the definition and then give the relevant background.
@@
We next define stability, which is a natural condition that rules out many pathological locally finite graphs, defined in [MR23, Section 4].
@@
    If $X$ is a tree with countable end space, then $X$ is stable [FGM21, Remark 5.5].
@@
    Non-stable locally finite graphs are more complicated and it will be easier to give examples after Definition 3.1. Intuitively, if there is a sequence of ends $\{\nu_{n}\}_{n}$ converging to an end $\mu$, none of which are dominated by $\mu$ and which are all of different types, then $\mu$ must not be stable. For concrete examples, see Example 4.22 and the first bullet point under part three in Definition 4.23. There are other ways to violate stability as well; the graphs in [MR24, Theorem 1.1] are not stable and contain a unique maximal end type.
@@
Mann and Rafi [MR23, Section 4] show that, without loss of generality, one can take $n=2$.
@@ 3.1 Signatures
By construction, signatures are collections of symbols associated to particular locally finite graphs which are invariant under proper homotopy. Throughout the paper, we sometimes mention signatures and actually mean a particular graph with that signature; for example, we may refer to the graph $(\omega^{3}+1)\vee C$. Also note that a proper homotopy equivalence of graphs induces a homeomorphism of their end spaces [AB02, Section 2]. Thus, for example, by an abuse of notation, we may refer to the end space of any graph with signature $(\omega^{3}+1)\vee C$ by $\partial((\omega^{3}+1)\vee C)$.
In the context of $(\{Y_{n}\}_{n},a)\to\mathcal{C}$, where each of the $Y_{n}$ are arbitrary locally finite trees, $\mathcal{C}$ is an infinite binary splitting tree, and $a$ is a specific bijection, a notation that is similar in flavor capturing behavior of the end space was developed by [KET78, Definition 0.4].
@@ 3.2 Building blocks
We now turn our attention to proving lemmas which will be useful throughout the paper. The following lemma is an implication of [MR23, Lemma 4.17], and we include a proof for completeness.
@@
First assume that $V$ is contained within $U$. It follows by taking $x=y$ in [MR23, Lemma 4.17] that all sufficiently small neighborhoods of $\nu$ are homeomorphic to $U$. In particular, if $W$ is a sufficiently small neighborhood of $\nu$, then there exists a homeomorphism $f:U\rightarrow W$ fixing $\nu$ and a homeomorphic copy $f(V)$ of $V$ in $W$. The set $f(V)$, being a subset of $W$, is also a sufficiently small neighborhood of $\nu$, and is therefore homeomorphic to $U$. Thus $U$ is homeomorphic to $V$. For the moreover statement, suppose that $\nu^{\prime}\sim\nu$. By [MR23, Lemma 4.17], all sufficiently small neighborhoods of $\nu^{\prime}$ are homeomorphic to $U$ via a homeomorphism taking $\nu$ to $\nu^{\prime}$, which immediately gives the result. ∎
@@
We first prove the statement when $A=\{\nu\}$ is a singleton. Without loss of generality, assume that $X$ is in the standard form of [AB02, Section 2]. Let $x_{0}$ be a vertex in $X$, and consider a non-backtracking path based at $x_{0}$ in $X$ limiting to $\nu$ with vertices $\{x_{n}\}_{n}$. Because $X$ is in standard form, $X\cong\{Y_{n}\}_{n}\rightarrow 1$, where each $Y_{n}$ is attached to the path limiting to $\nu$ via $x_{n}$. For each $n$, the set $\{\nu\}\cup(\bigcup_{i\geq n}\partial Y_{i})$ is clopen in $\partial X$. If, for contradiction, $V$ intersected infinitely many of these sets, then $\nu$ would be a limit point of $V$, contradicting that $V$ is closed. Thus take $U$ to be one of these sets for suitably large $n$.
@@
The end space $\partial X$ is compact and Hausdorff. By [AB02, Definition 2.5], it embeds into Cantor space, and therefore has a countable basis consisting of clopen sets. If $\partial X$ is not homeomorphic to Cantor space, then by Brouwer’s characterization of Cantor space, it has an isolated point. Hence $X$ has a discrete end. ∎
@@
Using the construction in [AB02, Definition 2.5], we fix an embedding of $\partial X$ into $\mathbb{R}$. For any clopen subset of the end space $U\subseteq\partial X$, there exist finitely many disjoint open intervals $I_{k}$ in $\mathbb{R}$ such that $\partial X\cap(\bigcup_{k}I_{k})=U$. Indeed, by Lemma 3.6 and compactness of $U$, there are finitely many open intervals in $\mathbb{R}$ covering $U$ with $\partial X\cap(\bigcup_{k}I_{k})=U$, and we may assume that they are disjoint because if two overlapped, we could combine them into one open interval. Thus for each $m$ we obtain finitely many disjoint open intervals $I_{m,1},...,I_{m,k_{m}}$ covering $U_{m}^{\prime}$ that are disjoint from $\partial X\setminus U_{m}^{\prime}$. $\{I_{m,i}\,:\,m,i\in\mathbb{Z}_{>0}$ and $1\leq i\leq k_{m}\}$ is a set of pairwise disjoint open intervals in $\mathbb{R}$, and thus we assume that they are indexed in an order-preserving way by $\mathbb{Z}$. Define $U_{m}:=I_{m}\cap\partial X$ so that $\{U_{m}\}_{m\in\mathbb{Z}}$ is a clopen decomposition of $\partial X\setminus V$.
@@ 4.1 Classification of local structures
Proposition 4.4 suggests that stability is a local version of self-similarity. Indeed, the bi-implication $1\Leftrightarrow 2$ is saying that a neighborhood of an end is stable if and only if a locally finite graph with end space equal to that neighborhood has self-similar end space. There is a construction in [MR24] of graphs with non-stable end spaces which contain a unique maximal end type such that this end type is of size one or infinity. Thus, the requirement in statement $3$ that maximal ends are stable is crucial. The graphs in [MR24] are not self-similar, and do not satisfy any of the conditions in Proposition 4.4.
@@ 4.2 The poset of local structures
Suppose for contradiction that the $\partial Z_{i}$ form a strictly decreasing sequence of local structures. We will inductively define a descending sequence $\{U_{i}\}_{i}$ of subsets of $\mathbb{R}$. We can homeomorphically map $\partial Z_{0}$ to some subset $U_{0}\subset\mathbb{R}$ of the real numbers by the construction in [AB02, Definition 2.5], and moreover, we may assume that the diameter of $U_{0}$, which we refer to as $\text{diam}(U_{0})$, is less than one in $\mathbb{R}$. Now suppose we have defined $U_{0},...,U_{i-1}$. Because $\partial Z_{i}\prec\partial Z_{i-1}$, there exists $\nu_{i}\in U_{i-1}$ such that $\nu_{i}\sim\partial Z_{i}$. We can realize a homeomorphic copy of $\partial Z_{i}$ as a subset $U_{i}$ of $U_{i-1}$ which is clopen in $U_{i-1}$ and which contains $\nu_{i}$. Furthermore, by stability, we may assume that $\text{diam}(U_{i})\leq\frac{1}{i+1}$. Thus, the $U_{i}$ are descending, and they are all compact and non-empty, with diameter approaching zero. This implies that $\bigcap_{i}U_{i}$ consists of a single point $x\in\mathbb{R}$. Let $\nu$ be the end in $\partial Z_{0}$ which maps to $x$, and let $\{V_{i}\}_{i}$ be a descending sequence of clopen subsets of $\partial Z_{0}$ such that $V_{i}\cong U_{i}$ for all $i\in\mathbb{Z}_{\geq 0}$. Note that as $Z_{0}$ is stable, the end $\nu$ must be stable too.